OpenStudy (stephanieelizzz):

Solve for x: log(1/3)+log(x)=log(6x-5) I got this to -->log(1/3x)=log(6x-5) but I'm not sure where to go from there.

OpenStudy (mathstudent55):

You did the first step correctly. Now use this: If \(\log x = \log y\), then \(x = y\).

OpenStudy (stephanieelizzz):

Not sure I get it. Would you be removing the logs since their bases are the same? @mathstudent55

OpenStudy (mathstudent55):

Correct. Using the same base log, if log A = log B then A = B. You have log(1/3 x) = log(6x - 5), then it must follow that 1/3 x = 6x - 5

OpenStudy (stephanieelizzz):

Gotcha! For some reason I'm ending up with a completely different answer(5/6.3) from what I'm supposed to be getting (15/17). Maybe I'm solving it incorrectly? @mathstudent55

OpenStudy (mathstudent55):

We'll do it together.

OpenStudy (mathstudent55):

\(\dfrac{1}{3}x = 6x - 5\) Multiply both sides by 3 to get rid of the denominator.

OpenStudy (mathstudent55):

\(x = 18x - 15\) Subtract 18x from both sides. \(-17x = -15\) Divide both sides by -17 \(x = \dfrac{15}{17} \)

OpenStudy (stephanieelizzz):

You're an angel. Thank you!!!

OpenStudy (mathstudent55):

You're welcome.